What is the 2001st number in the sequence: 2, 5, 8, 11,...?
(A) 5996 (B) 5999 (C) 6000 (D) 6001 (E) 6002

Found 2 solutions by manas, richwmiller:
Answer by manas(5)

answer is 6002

Answer by richwmiller(17219)

3n-1
3*(1)-1=2
3*(2)-1=5
3*3-1=8
3*4-1=11
3*2001-1=6003-1=6002